The temperature of a sample of copper increased by 23.0 degrees celsius when 255 J of heat was applied. What is the mass of the sampl?
Answer #1

heat applied = 255 Joule

temperature increased/difference = 23 oC

specific heat for copper = 0.385 J/g-oC

heat absorbed by sample of copper = mass * specific heat of copper * temperature difference

   255 Joule = mass * 0.385J/g-oC * 23o C

or mass = 255/(0.385*23) g = 28.797 grams